OpenStudy (anonymous):

A culture started off with 5000 bacteria after 5 hours it grew too 6500 bacteria predict how many bacteria will be present after 16 hours P=Ae^kt formula

OpenStudy (taramgrant0543664):

So you have to find k first so it started with 5000 so that's A P would be 6500 and t is 5 and you solve for k 6500=5000e^k(5)

OpenStudy (taramgrant0543664):

Once you find the k value then you can use that as k, use 5000 as A and your t would be 16 and you would solve for P

OpenStudy (anonymous):

Well how do I find out what k is?

OpenStudy (taramgrant0543664):

6500=5000e^5k 6500/5000=e^5k ln(6500/5000)=5k ln(6500/5000)/5=k

OpenStudy (anonymous):

So k would be 1.3?

OpenStudy (anonymous):

Or 0.26?

OpenStudy (taramgrant0543664):

No k=0.052 you didn't do the ln of it or divide by 5

OpenStudy (taramgrant0543664):

(ln1.3)/5=0.052

OpenStudy (taramgrant0543664):

ln is the inverse of e

OpenStudy (anonymous):

So my final equation would be p=5000e^0.052*16 ?

OpenStudy (taramgrant0543664):

Yes and solving for p is the answer you want

OpenStudy (anonymous):

Ok thanks you!(:

OpenStudy (taramgrant0543664):

No problem happy to help!

OpenStudy (anonymous):

So it seems as if my answer is wrong, what should I get for the answer?

OpenStudy (taramgrant0543664):

What did you get?

OpenStudy (anonymous):

1195.48, I don't see how I got this

OpenStudy (taramgrant0543664):

I got 11489.6

OpenStudy (anonymous):

Alright, I guess I'll just have to keep trying, thanks for the help though(: